2) The gene consists of 3 identical exons and 4 identical introns of the regions, and the introns consist of 120 nucleotides each, and the entire gene has 1470 nucleotides. How many codons will the pre-mRNA, each exon, mRNA have?


1
Expert's answer
2021-02-15T12:15:45-0500

We have entire gene of 1470 nucleotide. In this, there are 3 identical exons (assume E1,E2,E3=E) and 4 identical introns (I1, I2, I3, I4=I)

Now as mentioned, Introns, I=120 nucleotide in length

Thus 4 introns will be of 4*120=480 nucleotide in length


Total length of the gene is 1470, we know entire gene length=Total introns length + Total Exons length

or 1470=480+ Total Exons Length

or Total exons length=1470-480

=990 nucleotides

Now, this is for 3 exons, so for 1 exon the length is= 990/3

= 330 nucleotide


Again, we know that 1 codon corresponds to 3 bases or nucleotides

So, 330 nucleotides (since only exons are coding region only, thus corresponds to codons) will contain,

330/3=110 codons


Thus, 110 codons in the pre-mRNA, each exon, mRNA have.